OpenStudy (anonymous):

A man when asked how many hens and buffaloes he has, he told that his animals have 120 eyes and 180 legs. How many hens and buffaloes does he have?

OpenStudy (ajprincess):

Let the no of hens be x and no of buffaloes be y 2x+2y=120 x+y=60-(1) 2x+4y=180 x+2y=90-(2) (2)-(1)=>y=30 (1)=>x=60-30 =30

OpenStudy (anonymous):

assuming you don't have any mutated hens or buffaloes with 3 eyes or 4 legs, you can set up a system like this to solve...: